Al will rent a car for the weekend. He can choose one of two payment plans. The first plan costs $51.98 for two days plus 15 cents per mile. The second plan costs $39.98 for two days plus 20 cents per mile. How many miles does Al need to drive for the two plans to cost the same?

Answer by Alan3354(30993)

Al will rent a car for the weekend. He can choose one of two payment plans. The first plan costs $51.98 for two days plus 15 cents per mile. The second plan costs $39.98 for two days plus 20 cents per mile. How many miles does Al need to drive for the two plans to cost the same?
-----------------------
5198 + 15m = 3998 + 20m (m = miles)
1200 = 5m
m = 240 miles